Two similar coils which are separated by 2m are having radius of 1m and 80 turns have a common axis.Calculate the strength of magnetic field at a point midway between them on their common axis when a current is 0.2A .
Ans 0.04√2T
1
Expert's answer
2020-11-12T09:36:50-0500

The magnetic field along the axis of a ring z=1 m away from its center produced by a single coil is


"b=\frac{\mu_0}{2}\frac{NIr^2}{(z^2+r^2)^{\frac{3}{2}}}."

For two coils, it will be


"B=\mu_0\frac{NIr^2}{(z^2+r^2)^{\frac{3}{2}}}=7.109\space\mu\text{T}."